If a force of 80N extends a spring of natural length 8m by 0.4m what will be the length of the spring when the applied force is 100N
1
Expert's answer
2012-05-29T07:47:25-0400
delta(l) = k * F
0.4 = k * 80
k = 0.4 / 80 = 0.005 m/N
So, l = 8 + k *
100 = 8 + 0.005 * 100 = 8.5 m.